OpenStudy (anonymous):

A waitress fills bowls with soup. She puts 2/3 of a cup of soup in each bowl. How many bowls can the waitress prepare with 40 cups of soup?

OpenStudy (anonymous):

40 out of 2/3

OpenStudy (anonymous):

your job is to divide \(40\) by \(\frac{2}{3}\)

OpenStudy (anonymous):

\[40\div \frac{2}{3}=40\times \frac{3}{2}=20\times 3\]

OpenStudy (anonymous):

It is 60